INAF queue observing, 2016 Jan 24-25

Observer: F. Cusano, A. Rossi
Telescope Operator: David Gonzalo-Huerta
Support Astronomer: Barry Rothberg
Instrument: LUCI

Summary

At the sunset the sky is clear. Wind at the limit (17m/s)
Note: We can not observe the Iovino program because the mask was not loaded on the LUCi cabinet also if it was requested.

Night duration (12\xB0 el twilight) 19:20-6:10 ( 02:20-13:10 UT) = 10 h 50m

Observing time: * 7 h * *30 m (72%)

Weather time loss: 2 h (20%)

technical time loss 0.85 h (8%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Darks     10x60,3x300 good for Guaita, Gilli and DeBarros
Antoniucci EZ_TAU_E 200 H+K grating, zJspec, HKspec 5min completed
Antoniucci V1118 200 H+K grating,zJspec, HKspec 0.3h completed
Antoniucci V1647 200 H+K grating,zJspec, HKspec 0.3h completed
Guaita M0717 Br_gam 0.5h  
Gilli ID3605 200 H+K grating,zJspec 2.5h  
DeBarros MACS1423 200 H+K few min mask not centered, data not useful

Details

00:15 We start dark with LUCI
  • Guaita dark luci2.20160125.0007-0016
  • Gilli e De Barros dark luci2.20160125.0017-0021 wrong the camera was setted to N3.75 do again
  • Gilli e De Barros dark luci2.20160125.0022-0024

01:16 The TO opens the dome

01:45 slew to photometric standard of Guaita
  • seeing is too bad 2" we move to Antoniucci program

01:55 slew to Antoniucci program object UZ_TAU_E
  • acquisition of telluric, problem with LUCI, the mask is not moving (2:03)
  • seeing 2.2"

2:28 problem solved, acquisition EZ_TAU_E_tell.xml
  • seeing 2.2"
  • aquisition luci2.20160125.0025-0027
  • spectra luci2.20160125.0028-0031

2:39 slewing to EZ_TAU_E
  • acquisition luci2.20160125.0032-0036
  • spectra luci2.20160125.0037-0043
  • run imaging of EZ_TAU_E luci2.20160125.0044-0047, the grism is still inside but we can see the images, we stop the script and restart the scheduler, now is working luci2.20160125.0048-0053

3:14 slew to V1118 ORI
  • seeing 2.5"
  • imaging luci2.20160125.0054-0059
  • acquisition luci2.20160125.0060-0063
  • spectra luci2.20160125.0064-0071
  • seeing 1.8" at 3:42
  • seeing 1.4" at 3:47
  • telluric of V1118 Ori, acquisition luci2.20160125.0072-0074
  • spectra of telluric V1118 Ori luci2.20160125.0075-0079

4:06 slewing to V1647
  • imaging luci2.20160125.0080-0084
  • seeing 2.2" at 4:10
  • acquisition luci2.20160125.0085-0088
  • spectra luci2.20160125.0089-0096
  • seeing 1.3" at 4:33
  • seeing 1.2" at 4:40
  • seeing 1" at 4:57 seeing is improving we try to move to Guaita imaging program

5:00 slew to Guaita_BRgam
  • seeing 1.3 on the guider; we measure the seeing on the images and is 1", we keep observing
  • luci2.20160125.0097-126
  • seeing 1.1" at 5:12
  • some clouds at 5:25
  • seeing 1.4" at 5:48

5:50 slew to standard P161-D for Guaita Br_gamma and H2 filters luci2.20160125.0127-146
  • seeing 1.2" at 05:55 on image
  • seeing 1.3" at 6:05 on image
  • seeing 0.7" at 6:10 on image

6:14 slew to Guaita_H2 luci2.20160125.0147
  • 6:15 lost guide star

6:32 we try to change to Gilli,
  • telluric acquisition luci2.20160125.0148-0151
  • seeing 1.1" at 6:50
  • telluric spectra luci2.20160125.0152-00154

7:03 slew to Gilli science
  • acquisition luci2.20160125.0155-0158
  • seeing on the guider 1.3" at 7:05
  • spectra luci2.20160125.0159-0188
  • seeing on the guider 1" at 7:18
  • seeing on the guider 1.2" at 7:40
  • seeing on the guider 0.9" at 8:00
  • seeing on the guider 0.8" at 8:35
  • seeing 0.9" at 9:42
  • some clouds in the last exposure 188

10:02 we do again the telluric of Gilli because the airmass changed a lot
  • some clouds coming in
  • checking pointing
  • acquisition luci2.20160125.0189-0191
  • the sky is 80% covered by clouds at 10:08
  • spectra luci2.20160125.0192-0194

10:19 we slew to DeBarros telluric
  • acquisition luci2.20160125.0195-0198
  • need to check pointing 10:21
  • sky 60% covered by clouds, but it is free in the direction of the pointing
  • seeing 0.9" at 10:25
  • spectra luci2.20160125.0199-0206

10:41 moving to DeBarros science field
  • acquisition luci2.20160125.0207-0208
  • spectra luci2.20160125.0209-
  • seeing 1" at 11:00, some clouds
  • at 11:06 the cloud cover also part of the science field, guise star has fainted of 0.5 mag, some scattered light from the moon the collimation is not good
  • at 11:15 we lost the guide star

12:40 still cloudy, move to calibrations

13:00 calibration Gilli and Antoniucci zJspec
  • flat Halo2 luci2.20160125.215-219, 10sec
  • arc lamps see pdf file. sequences of lamp in, and 2 without lamp
  • arc lamp Argon luci2.20160125.220-222, Neon luci2.20160125.223-225, Xenon luci2.20160125.226-228, 10sec

13:20 calibration Antoniucci HKspec
  • luci2.20160125.229-235 too bright or too faint do not use! with HK Halo2 gives too many counts, switch to Halo3
  • flat Halo3 luci2.20160125.238-42, 40sec
  • arc lamps see pdf file. sequences of lamp in, and 2 without lamp
  • arc lamp Argon luci2.20160125.245-247, Neon luci2.20160125.248-250, Xenon luci2.20160125.251-253 , 10sec

13:49 flats Brgamma for Guaita
  • flat halo3, 10 sec luci2.20160125.254 wrong camera do not use
  • flat halo3, 10 sec luci2.20160125.255-259

-- %USERSIG{FeliceCusano - AndreaRossi - 2016-01-24}%

Comments

 
Topic revision: r16 - 17 Feb 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback